OpenStudy (anonymous):

If r is a linear function with a slope that is not equal to zero, which of the following is always true of its inverse function? A) Its slope is the negative reciprocal of the slope of r. B) Its slope is the reciprocal of the slope of r. C) Its y-intercept is the negative reciprocal of the y-intercept of r. D) Its y-intercept is the reciprocal of the y-intercept of r.

OpenStudy (freckles):

let's find out... say we have the function: \[f(x)=mx+b\] can you find the inverse function?

OpenStudy (freckles):

first step: solve y=mx+b for x

OpenStudy (anonymous):

Ok. So that's going to be X=mY+B right?

OpenStudy (freckles):

ok you switch x and y around but do you know how to solve that for Y?

OpenStudy (anonymous):

umm i forgot how to solve it. can u pls explain me how to solve it?

OpenStudy (anonymous):

wait so M is going to that X side right? im not sure

OpenStudy (freckles):

x=my+b you want to get y by itself so we have to undo the operations around it notice that you have a plus b on the side with the my subtract b on both sides

OpenStudy (freckles):

and yes we will later undo the operation of multiplication of m to that y by dividing both sides by m

OpenStudy (anonymous):

ohh yeah. i remember now. so that;s going to be mx-b=y?

OpenStudy (freckles):

So finding the inverse function is a two step process for linear functions: You have x-b=my divide both sides by m

OpenStudy (freckles):

x=my+b we subtracted b on both sides x-b=my+b-b x-b=my+0 x-b=my now we have one step to undo that multiplication of m onto that y so we divide m on both sides

OpenStudy (anonymous):

ok i got that now. then what is next?

OpenStudy (freckles):

well what do you have as a result of dividing m on both sides

OpenStudy (anonymous):

x/m-b=y?

OpenStudy (freckles):

close you divide both sides by m \[\frac{x-b}{m}=\frac{my}{m} \\ \frac{x-b}{m}=y \\ y=\frac{x-b}{m} \\ y=\frac{x}{m}-\frac{b}{m} \\ y=\frac{1}{m}x-\frac{b}{m} \\ f^{-1}(x)=\frac{1}{m}x-\frac{b}{m} \\ \text{ and we have } \\ f(x)=mx+b\]

OpenStudy (freckles):

notice the slope of f is m and the y-intercept is b notice the slope of inverse of f is 1/m and the y-intercept is -b/m

OpenStudy (anonymous):

ohh yeah i got it now. understand.

OpenStudy (freckles):

so what is your answer then

OpenStudy (anonymous):

is it B?

OpenStudy (freckles):

boom yeah or whatever the kids are saying these days

OpenStudy (anonymous):

haha thanks @freckles

OpenStudy (freckles):

np yes m and 1/m are reciprocals

OpenStudy (anonymous):

Alright. i got it now. appreciate it @freckles